UT 20150404 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: WSeifert, APramskiy, JHeidt, DThompson (all remote from Tucson)
Observer AO-Team: SEsposito, EPinna, APuglisi (all remote from Tucson)
Telescope Operator: GBechetti
LBTO Support: TSargent (remote from Tucson), DMiller, JChristou (Training)

Plan for the night

As per agreement, this night will be converted from engineering to LUCI2 AOCom night:

- Continue AOCom

- Training LBTO personnel (DMiller, JChristou)

Summary

  • Arcetri people trained DMiller & JChristou, who successfully operated the AO system afterwards!!!

  • Tests for accuracy of offsets, large offsets

  • Nice AO-image

  • Lots of network issues (LUCI-SW frozen frequently)

Preparations

LUCI2/AGw attached back to the telescope, functional checkout of both components

Details

All times are UT
Files are luci2.20150404.NNNN.fits

07:45 Sunset, sky mostly covered with thin/thick cirrus, but may improve during the night

T = 3deg, Hum = 30%, low wind

  • Check of system

- GEIRS readout Ok, ROM LUCI2 OK, IM initialized, all OK except for calibration lamps

02:45 Arcetri-group is done with their checkout, Gino opens, cirrus less dense

03:20 Gino can not find a bright star, homing mirrors, homing hexapod + tertiary, finally found one + collimate

  • Accuracy of offsets

03:50 A45_R8_0029, R+I=8.1, dice5,10times 3sec, H-band, about 2mga absorption, GEIRS stopped unexpectedly,
#1-2, no star in fieldA45_R8_0045_spec.xml

Arcetri-people doing some tests

04:20 HR3854, R+I ~ 6, PA=0, preset failed, Gino checks pointing, preset again, failed again

04:29 HR3854, R+I ~ 6, PA=90 (different GS), failed again, check pointing on a new M5 star

04:34 HR3854, R+I ~ 6, PA=90, failed again, new check pointing, something wrong

05:00 TYC2526, R+I ~ 6, PA=180, #4-6, OK,

Check accuracy of transform: Pos of star, #7 x,y=678/1290, #8=650/626 (gives 2.9deg?)
Put in N30 LS, throughslit, slitview, #9, gives 0.6deg => sign-wrong

05:38 Preset again, MOS error when trying to put back the N30 LS back to storage
same problem as earlier this week (rotator angle out of tolerance)
Walter put the mask back to storage until Peter fixes the problem
=> at present no presets via LUCI possible => Arcetri people train Doug/Julian

Peter corrected and checked masks:

=> First 10 masks + N30 Sieb V, N30 LS, N30 LS neu I, N30 LS neu II can be used without restriction

08:10 Preset to HD128197, dice5,10times 3sec, H+FeII, #179, image not loaded in Aladin, corrected
resent preset, OK; preset AO failed, restarted OB execution panel, OK, #180, actuators died, restart LUCI-SW

08:58 Preset to HD128197, dice5,10times 3sec, H+FeII, LUCI2-SW hangs again, several times tonight,
reason unclear

Logged into obs4 on the mountain and started LUCI2-SW from there

09:52 Preset to HD128197, dice5,10times, 3sec, H+FeII, finally!!!
#180-229, time to execution 30min

Align N30: FM4 M1: +2550steps M2: -2150steps

10:51 NGC6543, Br_gam, #230 than stopped, tried to start NCPA corection, WFS-star to faint ie bin=2,
#230-246, time to execution 35min, 10*1min on source, 6*1min offset
Test for big offsets => successfull!!

11:34 NGC6826, Br_gam, 1st AO-offset sent from LUCI failed (was only about 3.5"), tried to preset again, LUCI-SW got stuck, resent preset, AO arbitrator failed, reset shell
restarted, LUCI-SW got again stuck....

Networkproblems all night

12:48 Closing up.
Topic revision: r6 - 04 Apr 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback